Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night Report Page - 20120924

Observers : Iair Arcavi, Yen-Chen Pan. 
Data reduction team : K. Maguire, M. Sullivan, A. Gal-Yam, O. Yaron

Some distant clouds at sunset, then clear. Could be photometric.
SOFI photometric standard: bd174708003
EFOSC spectroscopic standard: gd71 (possible problem on the red end)
SN2012ca [EFOSC: Ugriz] - Followup
Seeing ~1.0".
EFOSC --> SOFI
SN2012ca [SOFI:Rg] - Followup
Seeing ~0.9". Had to change rotator angle to get rid of the noise in the upper right quadrant. 
Hip091690 Observed before and after the target (observed twice after the target because of the noise problem).
*** Noise problem continues. TO had to go to the telescope to fix it. ***
LSQ12dwl [SOFI: Rg] - Followup
Seeing ~0.9". Hip106951 observed after the target.
SN2012ec [SOFI: JHK Bg Rg] - Followup
Seeing ~ 0.8"-1.1". Hip12852 observed before and after the target.
SOFI --> EFOSC
*** Technical problems with EFOSC focusing and then shutter. ***
LSQ12ezn [EFOSC: gm13] - Followup blue continuum, squeezed in at twilight.
Pretty noisy but still blue.
Comments